Grade 11Physical Chemistry

The vapour density of a mixture containing NO2 and N2O4 is 38.3 at 27℃. The moles of NO2 in 100 moles mixture is

Saikiran
8 Years agoGrade 11

1 Answer

Hitesh Yadav
8 Years ago
Now in the mixture of 100 moles NO2 and N2O4, let us suppose that the number of moles of NO2 is x and hence number of moles of N2O4 is (100-x).Molar mass of NO2 is 14 + 2 × 16 = 46 gMolar mass of N2O4 is 2 × 14 + 4× 16 = 92 gNow if x moles of NO2 are present in mixture, it`s mass will be molar mass times the number of moles.So mass of NO2 = 46xSimilarly, mass Of N2O4 is 92(100-x)From the given question, vapour density of the mixture is given which is 38.3.By using Relation Molar mass = 2 × Vapour density of mixture, we can find the molar mass of the mixture. Which comes out to be 2 × 38.3 = 76.6 gHence over all molar mass would beMass of NO2 +Mass of N2O4 = Mass of the mixturePutting all the values we have46x + 92(100-x) = 76.6 ×100Solving for x we get x = 33.47 mole which gives us the number of moles of NO2 in the mixture